### v1

# A Spatio-Temporal Machine Learning Model for Mortgage Credit Risk: Default Probabilities and Loan Portfolios

Abstract
We introduce a novel machine learning model for credit risk by combining tree-boosting with a latent spatio-temporal Gaussian process model accounting for frailty correlation. This allows for modeling non-linearities and interactions among predictor variables in a flexible data-driven manner and for accounting for spatio-temporal variation that is not explained by observable predictor variables. We also show how estimation and prediction can be done in a computationally efficient manner. In an application to a large U.S. mortgage credit risk data set, we find that both predictive default probabilities for individual loans and predictive loan portfolio loss distributions obtained with our novel approach are more accurate compared to conventional independent linear hazard models and also linear spatio-temporal models. Using interpretability tools for machine learning models, we find that the likely reasons for this outperformance are strong interaction and non-linear effects in the predictor variables and the presence of large spatio-temporal frailty effects.
1 Introduction
Accurately assessing credit default risk is a challenging and critical task for financial institutions, investors, regulators, and policy makers. Traditionally, linear models such as linear discriminant analysis, logistic regression, or linear discrete hazard models have been used to model default probabilities of individual loans (Altman, 1968; Zmijewski, 1984; Shumway, 2001). Linear hazard models have been extended to account for temporal correlation that cannot be captured by observable covariates (Duffie et al., 2009; Koopman et al., 2011). Such residual correlation is also called frailty correlation. For loans associated with a spatial location, for instance, mortgages and loans to small and medium-sized enterprises (SMEs), there is likely spatial dependence among loans that cannot be fully explained by observable predictor variables, and several linear models have been proposed to account for this spatial frailty correlation (Fernandes and Artes, 2016; Agosto et al., 2019; Calabrese et al., 2019; Babii et al., 2019; Calabrese and Crook, 2020; Medina-Olivares et al., 2022; Calabrese et al., 2024). Recently, linear spatio-temporal models for credit risk have been introduced which model space-time correlation but assume a linear functional form in the predictor variables (Berloco et al., 2023; Medina-Olivares et al., 2023b). However, non-linear machine learning models often achieve a higher prediction accuracy than linear models (Barboza et al., 2017; Zieba et al., 2016; Xia et al., 2017; Sigrist and Hirnschall, 2019; Sigrist and Leuenberger, 2023; Cheraghali and Molnár, 2024). To the best of our knowledge, there exists no prior work that uses state-of-the-art machine learning models and explicitly accounts for spatio-temporal frailty correlation for modeling credit default risk.
In this article, we introduce a novel approach which combines tree-boosting with a latent spatio-temporal Gaussian process. This allows for modeling non-linear and interaction effects of predictor variables as well as for accounting for spatio-temporal frailty correlation among loans which is not accounted for by the observable predictor variables. It is likely that not all relationships are linear, and more realistic non-linear models allow for gaining a better understanding of default mechanisms and for generating more accurate predictions. Furthermore, the space-time Gaussian process allows for generating spatially and temporally varying frailty default risk maps, which can provide valuable insights, while also improving overall prediction accuracy. We apply our proposed model to a large U.S. mortgage data set and compare it to a linear hazard, a linear spatial, and a linear spatio-temporal model regarding the accuracy of predictive default probabilities for individual loans and predictive loan portfolio loss distributions. We first observe that, compared to an independent linear hazard model, incorporating spatial frailty correlation improves the prediction accuracy of default probabilities. Additionally modeling space-time correlations further increases the prediction accuracy compared to a purely spatial linear model. Further, we find that the predictive default probabilities of our proposed spatio-temporal machine learning model are more accurate compared to the default probabilities of a linear spatio-temporal model. Concerning loan portfolios, we analyze the accuracy of predictive loan portfolio loss distributions using the continuous ranked probability score (CRPS) as well as the accuracy of 99% upper quantiles using the corresponding quantile loss. We find that our proposed spatio-temporal frailty machine learning model results in more accurate predictive loan portfolio loss distributions compared to all linear models considered, i.e., a linear hazard model, a linear spatial model, and a linear spatio-temporal model. Interestingly, when considering predictive means and predictive upper quantiles of loan portfolio distributions in the years of the global financial crisis around the year 2009, our tree-boosted spatio-temporal frailty model predicts a higher and thus more realistic loss at the beginning of the crisis, and the loss then reverts faster to lower and more accurate levels after the crisis compared to the other models considered. We also investigate the reasons for the higher prediction accuracy of our proposed spatio-temporal machine learning model, and we find two main explanations. First, there is considerable spatio-temporal variation that is not captured by observable predictor variables. Second, we find that the tree-boosting part of the proposed model accounts for interactions and non-linear effects in the predictor variables that cannot be captured with a linear functional form.
The remainder of this article is organized as follows. In Section 2 we introduce the methodology, and in Section 3, we apply and compare our methodology on a large U.S. mortgage data set. Section 4 concludes.
2 Spatio-temporal frailty correlation and tree-boosting
2.1 Notation and default probabilities
Our goal is to model default events of $N\in\mathbb{N}^{+}$ loans. For every loan $i$, $i=1,\dots,N$, we assume that we observe predictor variables $X_{ki}\in\mathbb{R}^{p}$ at discrete times $t_{ki}$, $k=0,\dots,n_{i}-1$, and a default time $\tau_{i}\in\{t_{1i},\dots,t_{n_{i}i},\infty\}$, where $0\leq t_{0i}\leq t_{ki}\leq t_{n_{i}i}\leq T$, $k=0,\dots,n_{i}$. I.e., $X_{ki}$ are the predictor variables of loan $i$ observed at time $t_{ki}$, and $\tau_{i}=t_{ki}>t_{0i}$ means that loan $i$ has defaulted in the interval $(t_{k-1i},t_{ki}]$. Further, $t_{0i}$ denotes the time when a loan $i$ enters the set of active loans, $t_{n_{i}i}$ denotes the last observation time for loan $i$, and $n_{i}$ is the total number of temporal observations for loan $i$. The last observation time $t_{n_{i}i}$ can be either the default time $\tau_{i}$, the time of some other form of exit such as reaching the loan maturity date, or the end of the observation period $T$. In addition, we assume that every loan $i$ has associated spatial coordinates $s_{i}\in\mathcal{D}\subset\mathbb{R}^{2}$.
Let $P\left(\tau_{i}=t_{k+1i}\middle|\tau_{i}>t_{ki}\right)$ denote the probability that a loan $i$ defaults in the interval $(t_{ki},t_{k+1i}]$ given that it has not defaulted until time $t_{ki}$. In a traditional independent linear hazard model, it is assumed that
$P\left(\tau_{i}=t_{k+1i}\middle|\tau_{i}>t_{ki}\right)=\left(1+e^{-X_{ki}^{T}\beta}\right)^{-1},\leavevmode\nobreak\ \leavevmode\nobreak\ \beta\in\mathbb{R}^{p}.$ (1)
Assuming independence across space and time conditional on $X_{ki}$, the corresponding likelihood is given by
$\prod_{i=1}^{N}\prod_{k=0}^{n_{i}-1}\left(\mathbbm{1}_{\{\tau_{i}=t_{k+1i}\}}\left(1+e^{-X_{ki}^{T}\beta}\right)^{-1}+\mathbbm{1}_{\{\tau_{i}>t_{k+1i}\}}\left(1-\left(1+e^{-X_{ki}^{T}\beta}\right)^{-1}\right)\right).$ (2)
2.2 Accounting for spatial and spatio-temporal frailty correlation
In the following, we first lift the assumption of independence between loan-time observations by introducing latent frailty variables that model spatial and spatio-temporal correlation. Instead of (1), we assume
$P\left(\tau_{i}=t_{k+1i}\middle|\tau_{i}>t_{ki},b(t_{ki},s_{i})\right)=f(F(X_{ki})+b(t_{ki},s_{i})),$ (3)
where $f(\cdot)$ is a link function such as $f(x)=\left(1+e^{-x}\right)^{-1}$, $F(\cdot)$ is a function $F:\mathbb{R}^{p}\mapsto\mathbb{R}$, and the latent variable $b(\cdot,\cdot)$ is a zero-mean Gaussian process (Williams and Rasmussen, 2006) that accounts for spatial or spatio-temporal frailty correlation. Both $b(\cdot,\cdot)$ and $F(\cdot)$ are specified in the following.
For the latent frailty process $b(t,s)$, we consider two cases: a spatial model and a spatio-temporal model. In the spatial model, the Gaussian process $b(t,s)$ varies over space only and is constant over time. It is defined by a spatial covariance function $\text{Cov}(b(t,s),b(t^{\prime},s^{\prime}))=c_{\theta}(s,s^{\prime})$, $s,s^{\prime}\in\mathcal{D}$, which depends on a set of parameters $\theta\in\Theta\subset\mathbb{R}^{q}$. For instance, in the application below, we use a Matérn covariance function
$c_{\theta}(s,s^{\prime})=\sigma_{1}^{2}\frac{2^{\nu-1}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}||s-s^{\prime}||}{\rho_{s}}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}||s-s^{\prime}||}{\rho_{s}}\right),$ (4)
where $\sigma_{1}^{2}$ is a marginal variance parameter, $\rho_{s}$ is a spatial range, or scale, parameter, $\nu$ is a smoothness parameter, $\Gamma(\cdot)$ is the Gamma function, and $K_{\nu}(\cdot)$ is the modified Bessel function of the second kind. For spatio-temporal models, the Gaussian process varies over both space and time, and it is defined by a space-time covariance function $\text{Cov}(b(t,s),b(t^{\prime},s^{\prime}))=c_{\theta}((t,s),(t^{\prime},s^{\prime}))$, $(t,s),(t^{\prime},s^{\prime})\in[0,T]\times\mathcal{D}\subset\mathbb{R}^{3}$. In our application, we use an anisotropic spatio-temporal Matérn covariance function:
$c_{\theta}((t,s),(t^{\prime},s^{\prime}))=\sigma_{1}^{2}\frac{2^{\nu-1}}{\Gamma(\nu)}\left(\sqrt{2\nu}||A^{-1}((t,s)-(t^{\prime},s^{\prime}))||\right)^{\nu}K_{\nu}\left(\sqrt{2\nu}||A^{-1}((t,s)-(t^{\prime},s^{\prime}))||\right),$ (5)
where $A=\text{diag}(\rho_{t},\rho_{s},\rho_{s})$ is a diagonal matrix containing a temporal and spatial range parameters $\rho_{t}$ and $\rho_{s}$, respectively.
The default probability $P\left(\tau_{i}=t_{k+1i}\middle|\tau_{i}>t_{ki},b(t_{ki},s_{i})\right)$ in (3) is conditional on $b(t_{ki},s_{i})$, and the joint marginal likelihood of all loans and time points does not simply factorize as in (2). It is given by
[…]
where $b$ and $F$ denote the stacked evaluation of $b(\cdot,\cdot)$ and $F(\cdot)$ at all data points, i.e.,
$\begin{split}b&=(b(t_{01},s_{1}),\dots,b(t_{n_{1}-11},s_{1}),b(t_{02},s_{2}),\dots,b(t_{n_{2}-12},s_{2}),\dots,b(t_{0N},s_{N}),\dots,b(t_{n_{N}-1N},s_{N}))^{T},\\ F&=(F(X_{01}),\dots,F(X_{n_{1}-11}),F(X_{02}),\dots,F(X_{n_{2}-12}),\dots,F(X_{0N}),\dots,F(X_{n_{N}-1N}))^{T},\end{split}$
$p(b|\theta)$ denotes the density of $b$, and $\mathcal{L}_{ki}(F(X_{ki}),b(t_{ki},s_{i}))$ is defined as
$\mathcal{L}_{ki}(F(X_{ki}),b(t_{ki},s_{i}))=\mathbbm{1}_{\{\tau_{i}=t_{k+1i}\}}f(F(X_{ki})+b(t_{ki},s_{i}))+\mathbbm{1}_{\{\tau_{i}>t_{k+1i}\}}\left(1-f(F(X_{ki})+b(t_{ki},s_{i}))\right).$
If $F(\cdot)$ is a linear function, $F(X_{ki})=X_{ki}^{T}\beta$, we refer to a model as defined in (3) as a “linear spatial model” when the latent Gaussian process varies over space only and as a “linear spatio-temporal model” when the Gaussian process varies over both space and time.
2.3 A tree-boosted spatio-temporal Gaussian process model
In the following, we show how to relax the linearity assumption for the fixed effects predictor variable function $F(\cdot)$ using tree-boosting in models with latent Gaussian processes. Tree-boosting (Friedman et al., 2000; Friedman, 2001; Bühlmann and Hothorn, 2007; Sigrist, 2021) is a machine learning technique that often achieves superior prediction accuracy on tabular data sets (Nielsen, 2016; Shwartz-Ziv and Armon, 2022; Januschowski et al., 2022; Grinsztajn et al., 2022). For instance, recently Grinsztajn et al. (2022) showed that tree-boosting outperforms random forest and various state-of-the-art deep neural networks on a large collection of data sets. Sigrist and Leuenberger (2023) and Cheraghali and Molnár (2024) find similar results for credit risk data.
We assume the model specified in (3) and the subsequent paragraph. In addition, we assume that $F(\cdot)$ is a function in a normed function space $\mathcal{H}$ that is the linear span of a set $\mathcal{S}$ of base learners $f_{j}(\cdot):\mathbb{R}^{p}\rightarrow\mathbb{R}$, which consist of regression trees (Breiman et al., 1984) in this article. Our goal is then to find a joint minimizer for $F(\cdot)\in\mathcal{H}$ and $\theta\in\Theta$ of the functional obtained when plugging $F(\cdot)$ and $\theta$ into the negative log-likelihood:
$(\hat{F}(\cdot),\hat{\theta})=\operatorname*{argmin}_{(F(\cdot),\theta)\in(\mathcal{H},\Theta)}-\log\left(\mathcal{L}(F,\theta)\right)\Big{|}_{F=F(X)},$ (7)
where $F(X)$ is the row-wise evaluation of $F(\cdot)$ at $X\in\mathbb{R}^{n\times p}$, which is the matrix containing predictor variables for all observations, $n=\sum_{i=1}^{N}n_{i}$, and $\mathcal{L}(F,\theta)$ is given in (6). The minimization of the empirical risk functional in (7) is done iteratively using the latent Gaussian model boosting (LaGaBoost) algorithm given in Algorithm 1, which performs a form of functional gradient descent. In detail, this algorithm iterates between, first, finding a maximum for $\theta$ of $\mathcal{L}(F_{m-1},\theta)$ conditional on the current estimate $F_{m-1}(\cdot)$ and, second, updating the ensemble of trees $F(\cdot)$ using one functional gradient descent step given the current estimates $\theta_{m}$ and $F_{m-1}(\cdot)$. Specifically, the boosting update $f_{m}(\cdot)$ in iteration $m$ is given by the least squares approximation to the vector obtained when evaluating the negative functional gradient of the functional defined in (7) at $(F_{m-1}(\cdot),I_{X_{ki}}(\cdot))$, where $I_{X_{ki}}(\cdot)$ are indicator functions which equal $1$ at $X_{ki}$ and 00 otherwise. Equivalently, $f_{m}(\cdot)$ is the minimizer of a first-order functional Taylor approximation of the functional in (7) with $F(\cdot)=F_{m-1}(\cdot)+f(\cdot)$ around $F_{m-1}(\cdot)$ with an $L^{2}$ penalty on $f(\cdot)$ evaluated at $(X_{ki})$. For more details on the LaGaBoost algorithm, e.g., the calculation of functional gradients; see Sigrist (2022, 2023). In the following, we refer to such a model as “tree-boosted spatio-temporal frailty model” or sometimes shortly as “spatio-temporal LaGaBoost model”.
Input : Initial values $\theta_{0}\in\Theta$, learning rate $\nu>0$, number of boosting iterations $M\in\mathbb{N}$
Output : Function $\hat{F}(\cdot)=F_{M}(\cdot)$ and parameters $\hat{\theta}=\theta_{M}$
1:  Initialize $F_{0}(\cdot)=\operatorname*{argmax}_{c\in\mathbb{R}}\mathcal{L}(c\cdot 1,\theta)$
2:  for $m=1$ to $M$ do
3:     Find $\theta_{m}=\underset{\theta\in\Theta}{\operatorname*{argmax}}\mathcal{L}(F_{m-1},\theta)$ using a method for convex optimization initialized with $\theta_{m-1}$
4:     Find $f_{m}(\cdot)=\underset{f(\cdot)\in\mathcal{S}}{\operatorname*{argmin}}\left\|\frac{\partial\log(\mathcal{L}(F_{m-1},\theta_{m}))}{\partial F}-f\right\|^{2}$
5:     Update $F_{m}(\cdot)=F_{m-1}(\cdot)+\nu f_{m}(\cdot)$
6:  end for
Algorithm 1 LaGaBoost: Latent Gaussian model Boosting
In order that we can do estimation and prediction in a computationally feasible manner on large data sets in practice, we need to apply some approximations for both the linear Gaussian process and the tree-boosted Gaussian process models. In the following, we describe these approximations. First, the integral in (6) cannot be calculated in closed form, and we approximate it using a Laplace approximation. Laplace approximations are computationally efficient, have asymptotic convergence guarantees, and are accurate for large data sets; see, e.g., Kündig and Sigrist (in press). Furthermore, to ensure that computations with Gaussian processes scale to large data sets, we use Vecchia approximations (Vecchia, 1988; Datta et al., 2016; Katzfuss and Guinness, 2021) for the latent Gaussian processes $b$. In spatial statistics, Vecchia approximations have recently “emerged as a leader among the sea of approximations” (Guinness, 2021) and are often considered as “the most promising class of approximations” (Kang and Katzfuss, 2023). In brief, Vecchia approximations generate an approximate sparse reverse Cholesky factor of the precision matrix of the latent Gaussian process $b$ by using an ordered conditional approximation for the density of the Gaussian process. In doing so, every row of this sparse Cholesky factor contains maximally $m$ non-zero entries corresponding to nearest neighbors for every observation. For prediction, a Vecchia approximation is applied to the joint distribution of a latent Gaussian process at the training and prediction points. Posterior predictive distributions for the latent Gaussian process are then obtained as conditional distributions of this approximated joint distribution. Further, predictive probabilities for the observable response variables are calculated by numerically integrating over the posterior predictive distribution of the latent Gaussian variable using, e.g., adaptive Gauss-Hermite quadrature (Liu and Pierce, 1994). See Sigrist (2023) and Kündig and Sigrist (in press) for more information on Vecchia-Laplace approximations including estimation and prediction. Additionally, for increased computational efficiency, we use the iterative methods of Kündig and Sigrist (in press) for estimation and prediction with Vecchia-Laplace approximations instead of Cholesky decompositions.
3 Application to mortgage credit risk data
In the following, we apply our proposed models to a large U.S. mortgage credit risk data set.
3.1 Data and default definition
We consider mortgage data from Freddie Mac’s publicly available single-family loan-level data set***https://www.freddiemac.com/research/datasets/sf-loanlevel-dataset from release 37. Freddie Mac provides monthly loan-level credit performance records on all mortgages that Freddie Mac purchased or guaranteed from 1999 onwards. In addition, Freddie Mac provides for every year random subsamples of 50’000 loans that originated in the corresponding year. For our analysis, we consider the union of all fully amortizing 30-year fixed-rate mortgages in all random subsets of Freddie Mac from 1999 through 2022. We model the data at a yearly frequency, and for every year, our data set contains all active mortgages that have not been terminated at the beginning of the year and that have originated before this year in one of the subsamples of Freddie Mac. A mortgage is considered as terminated when the loan defaults, reaches maturity, or when its balance is reduced to zero for example due to prepayment or sale to a third party. Furthermore, a loan is considered to be in default if it is at least 90 days delinquent. For each year and all active mortgages, we construct a default indicator that equals one if a mortgage defaults during the year and zero otherwise. We do not consider a mortgage to be active at the beginning of the year if the most recent monthly performance record is older than six months or if no performance record is available for the first three months after the loan’s origination date. In addition, we restrict our sample to loans on properties located in the contiguous United States.
In total, our data set contains 2’256’528 loan-year observations for 538’942 different mortgages, of which 35’923 loans defaulted. Figure 1 shows the number of defaults and the default rate over time. Many defaults occur around the year 2009 during the global financial crisis, and a particular one-year spike occurs in the year 2020 during the COVID-19 pandemic.
Figure 1: Number of defaults and default rate over time.
For the single-family loan-level data set, Freddie Mac provides the first three digits of a five-digit postal code for the property of each mortgage. For confidentiality reasons, exact locations are not available. Every three-digit postal code is associated with a specific area in the United States, and we assign the centroid coordinates of the corresponding area to every mortgage. Our data set contains mortgages from a total of 875 different areas. Figure 2 shows the aggregate default rates for the three-digit postal code areas over the years 2000 to 2022. Spatial patterns are visible such as higher default rates in the states of California, Florida, Nevada, and Arizona, and lower default rates in states such as Idaho, Wyoming, North Dakota, and Nebraska.
Figure 2: Spatial default rates. No data is available for the gray areas.
3.2 Predictor variables
From the data provided by Freddie Mac, we construct several static and temporally varying predictor variables. The latter predictor variables are based on information available at the beginning of each year and include the age of the loan in months, the current loan-to-value ratio, and the interest rate spread. The current loan-to-value ratio is calculated by dividing the current unpaid principal balance by the appraised value of the property at the time of origination. To calculate the interest rate spread, we follow Calabrese et al. (2024) and subtract from the loan interest rates the average interest rate currently issued by Freddie Mac for 30-year fixed-rate mortgages. This average interest rate is retrieved from the Federal Reserve Bank of St. Louishttps://fred.stlouisfed.org/series/MORTGAGE30US. In addition, we use year-fixed effects. Static predictor variables include the borrower’s credit score at the time when the mortgage originated, additional financial indicators, and several characteristics about the property and the mortgage. A description of all predictor variables is provided in Table 1. These predictor variables are commonly chosen when Freddie Mac’s loan-level data set is used in the credit risk literature; see, e.g., Hu and Zhou (2019) and Medina-Olivares et al. (2023a). Missing values for categorical predictor variables are imputed using the most frequent category, and numeric predictor variables are imputed using the mean. Summary statistics of the numeric and categorical variables are provided in Tables 5 and 6, respectively, in Appendix A.5. For all linear models, we additionally include an intercept term in the predictor variables $X_{ki}$.
Variable Description
credit_score Borrower’s credit score
longitude Longitude of the centroid
latitude Latitude of the centroid
occupancy Indicates whether the property is owner-occupied (P), a second home (S), or an investment property (I).
nr_units Indicates whether the property has 1, 2, 3, or 4 units.
loan_purpose Indicates whether the mortgage is for cash-out refinance (C), no cash-out refinance (N), or for purchase (P).
first_time_homebuyer Indicates whether the borrower has not owned any residential property in the three years prior to the purchase of the mortgaged property.
msa Indicates whether the property is reported to be located in a metropolitan statistical area or not.
insurance_percent Percentage of loss coverage on the loan that an insurer is providing in the event of a loan default.
orig_dti Original dept-to-income: borrower’s monthly debt payments divided by the borrower’s monthly income
orig_cltv Original combined loan-to-value: original mortgage loan plus any secondary mortgage loan divided by the mortgage appraised value
orig_upb Original unpaid principal balance of the mortgage
multiple_borrowers Indicates whether more than one borrower is obligated to repay.
year_versioning Year of the loan-year observation
cnt_ltv Current loan-to-value: current unpaid principal balance divided by the mortgage appraised value
ir_spread Interest rate spread
n_months Age of the loan in months
Table 1: Description of predictor variables.
3.3 Models considered and implementation details
We consider the following models with increasing levels of complexity: an independent linear hazard model, a linear spatial model, a linear spatio-temporal model, and a tree-boosted spatio-temporal frailty model; see Section 2. For the spatial and spatio-temporal Gaussian process models, we use a Matérn covariance function as defined in Equations (4) and (5) with smoothness parameter $\nu=1.5$ and $m=20$ nearest neighbors for estimation and prediction with Vecchia approximations. For purely spatial processes, we use the Euclidean distance to determine the nearest neighbors in Vecchia approximations and a random ordering of the spatial coordinates since this gives accurate approximations (Guinness, 2018). For the spatio-temporal models, we use a correlation-based approach to determine nearest neighbors as in Kang and Katzfuss (2023) and an increasing temporal order and a random spatial order for coordinates with the same time. Following Kang and Katzfuss (2023), the nearest neighbors for the spatio-temporal models are redetermined in every iteration that is a power of two whenever an optimization algorithm is used for determining the parameters $\theta$. For prediction with Vecchia approximations, the observed points appear first in the ordering, and the Gaussian process at prediction points is only conditioned on the training data in the Vecchia approximation. For finding optima for the parameters $\theta$, we use the limited-memory BFGS algorithm for the linear Gaussian process models and gradient descent with Nesterov acceleration for the tree-boosted spatio-temporal frailty model. Estimation and prediction with the linear spatial, linear spatio-temporal, and the tree-boosted spatio-temporal frailty model is done using the GPBoosthttps://github.com/fabsig/GPBoost library version 1.4.0 (Sigrist et al., 2021). The code for executing the mortgage credit risk application and generating the data set is publicly available; see https://github.com/pkuendig/SpaceTimeFrailty.
3.4 Sample split for model evaluation and choosing tuning parameters
We conduct one-year-ahead default predictions for each year starting from 2008 through 2022 using an expanding window training data approach. I.e., we first consider all loan-year observations up to and including the year 2007 as training data and make predictions for all loan-year observations of 2008. We then continue by expanding the training window by one year and using the subsequent year as test data. I.e., all the following results are based on temporal out-of-sample predictions.
For choosing tuning parameters, we analogously split every training data set into an inner training data set and a validation data set. The validation data contains all loan-year observations of the most recent year in the training data, and the inner training data consists of all samples excluding this most recent year. Tuning parameters are chosen by estimating models on the inner training data and selecting the combinations of tuning parameters that maximize the area under the receiver operating characteristic curve (AUC) on the validation data. The candidate tuning parameters for the tree-boosted spatio-temporal frailty model are shown in Table 4 in Appendix A.4.
3.5 Prediction of individual default probabilities
We first apply the above-described models to predict one-year-ahead default probabilities of individual mortgage loans for every year from 2008 through 2022. The predictive default probabilities are evaluated using the following prediction accuracy measures: the AUC, the H-measure, the average log-loss, the Brier score, and the expected calibration error (ECE). The AUC measures predictive discrimination ability and can be interpreted as the probability that the predictive probability for a randomly drawn default event is higher than the predictive probability for a randomly drawn loan-year observation without a default. The ECE assesses calibration. A predictive probability $p$ is calibrated if the corresponding event (=default) occurs in $100\times p$ percent. We determine the boundaries of the bins for the ECE by using $20$ equally-spaced quantiles of the empirical distribution of all predictive probabilities of all models and years pooled together. The H-measure, the log-loss, and the Brier score measure overall predictive accuracy. See, e.g., Dimitriadis et al. (2023) for more information on these prediction accuracy measures.
Table 2 reports the AUC, the H-measure, mean log-loss, Brier score, and ECE for every model, averaged over the 15 years for which we perform one-year-ahead default predictions. We find that the predictive default probabilities of the tree-boosted spatio-temporal frailty model are the most accurate for all metrics considered. An independent linear hazard model has overall the worst prediction accuracy. Adding a spatial frailty variable to an independent linear hazard model improves the prediction accuracy. Further, the linear spatio-temporal model outperforms a purely spatial linear model in all prediction accuracy measures. We interpret this in the sense that there are both non-linearities and/or interactions among the predictor variables as well as spatio-temporal frailty effects present in the mortgage data. We corroborate this conclusion when analyzing the estimated models in more detail in Section 3.7.
AUC H-measure Mean log-loss Brier score ECE
Linear independent 0.7649 0.1951 $8.101\times 10^{-2}$ $1.682\times 10^{-2}$ $1.177\times 10^{-2}$
Linear spatial 0.7664 0.1981 $8.092\times 10^{-2}$ $1.686\times 10^{-2}$ $1.188\times 10^{-2}$
Linear spatio-temporal 0.7711 0.2079 $8.032\times 10^{-2}$ $1.679\times 10^{-2}$ $1.164\times 10^{-2}$
LaGaBoost spatio-temporal 0.7769 0.2148 $\mathbf{7.859\times 10^{-2}}$ $\mathbf{1.638\times 10^{-2}}$ $\mathbf{1.017\times 10^{-2}}$
Table 2: AUC, H-measure, average log-loss, Brier score, and ECE averaged over the 15 years for which one-year-ahead default predictions are calculated.
Figure 3 additionally reports the AUC of every model over time. We observe that the tree-boosted spatio-temporal frailty model has the highest AUC for most years. Further, the AUC is at a relatively high level for all models and for all years before the COVID-19 pandemic. In particular, the AUC remains at constant high levels during the global financial crisis around the year 2009. However, in the year 2020 during the COVID-19 pandemic, the AUC drops to considerably lower levels for all models and increases again in the subsequent years. Similar patterns over time are observed for the H-measure and the ECE; see Figures 10 and 11 in Appendix A.1. We interpret this sudden decrease in prediction accuracy of all models in 2020 in the sense that an artificial, external shock in the form of lock-downs led to very different default mechanisms compared to normal times and also compared to the global financial crisis around the year 2009.
Figure 3: Temporal out-of-sample test area under the receiver operator curve (AUC) (higher = better).
3.6 Prediction of loan portfolio loss distributions
In the following, we apply the different models for predicting one-year-ahead loss distributions of annual mortgage portfolios containing all active loans from the beginning of every year. If a loan defaults, this results in a loss corresponding to the loan’s unpaid principal balance at the time when predictions are made. Predictive loss distributions of portfolios are approximated by simulating 100’000 times sums of Bernoulli variables with according predictive default probabilities. For the models with a latent Gaussian process, we use a two-step simulation approach as follows. In every simulation run, first, a sample from the posterior predictive distribution of the latent Gaussian process is drawn which is then added to the predictions of the fixed effects function to obtain predictive probabilities which, in a second step, are used to simulate default indicator variables.
[…]
Table 3 reports the CRPS, the 99% quantile loss, and the RMSE for the 15 years used as test data. We observe that the tree-boosted spatio-temporal model has the lowest CRPS, quantile loss, and RMSE compared to all other models. I.e., predictive portfolio loss distributions of the tree-boosted spatio-temporal model are the most accurate when looking at the entire distribution, upper quantiles, and the center of the distribution. Further, the predictive portfolio loss distributions of the linear spatio-temporal model are more accurate in terms of the CRPS and the 99% quantile loss compared to the linear spatial model and also the independent linear hazard model. The linear spatial model has a lower 99% quantile loss but a higher RMSE and CRPS than the independent linear model. I.e., as expected, adding a spatial or a spatio-temporal frailty variable improves the accuracy of upper tail predictions compared to a linear model assuming independence conditional on observable predictor variables.
CRPS 99% quantile loss RMSE
Linear independent $2.873\times 10^{8}$ $1.237\times 10^{8}$ $5.082\times 10^{8}$
Linear spatial $2.882\times 10^{8}$ $1.229\times 10^{8}$ $5.088\times 10^{8}$
Linear spatio-temporal $2.790\times 10^{8}$ $1.125\times 10^{8}$ $5.148\times 10^{8}$
LaGaBoost spatio-temporal $\mathbf{2.553\times 10^{8}}$ $\mathbf{1.058\times 10^{8}}$ $\mathbf{4.710\times 10^{8}}$
Table 3: Accuracy of one-year-ahead predictive loan portfolio loss distributions.
Next, Figure 4 shows the differences between the means of the predictive portfolio loss distributions and the realized portfolio losses over the years 2008 to 2014. Interestingly, among the models considered, the tree-boosted spatio-temporal frailty model predicts the highest and most realistic mean portfolio loss at the beginning of the global financial crisis during the years 2008 and 2009, and subsequently, its mean portfolio loss decreases faster to more realistic levels after the crisis. From a risk management perspective, such a lower pre-crises underestimation and lower post-crisis overestimation of portfolio loss is clearly desirable. In Figure 12 in Appendix A.2, we also show the differences between the predictive mean portfolio losses and the realized portfolio losses for all prediction years from 2008 to 2022.
Figure 4: Differences between means of the predictive loss distributions and realized portfolio losses.
Figure 5 additionally shows the time series of predictive $99\%$ quantiles of one-year-ahead portfolio loss distributions. Similar to the mean portfolio loss, we find that, among the models considered, the tree-boosted spatio-temporal model predicts the highest upper tail loan portfolio loss at the beginning of the global financial crisis and its predictive upper tail losses are the smallest after the crisis. The independent model without a frailty Gaussian process generates the smallest predictive $99\%$ quantiles before the global financial crisis. The sudden one-year default spike in 2020 during the COVID-19 pandemic causes all classifiers to predict high upper tail loan portfolio losses for the following year 2021.
Figure 5: Predictive $99\%$ quantiles of one-year-ahead loan portfolio loss distributions.
3.7 Model interpretation
In the following, we aim to better understand the functioning of the tree-boosted and linear spatio-temporal frailty models. In Figure 6, we show the mean of the posterior distribution of the latent spatio-temporal frailty Gaussian process for the tree-boosted model when training on data up to and including the year 2021, which corresponds to the model for predicting defaults in the most recent year 2022. Due to space constraints, posterior means are not shown for the year 2000. We additionally show these posterior mean maps for the linear spatio-temporal model in Figure 13 in Appendix A.3. We observe that there is considerable variation in the latent frailty variable over space and this variation changes over time. In particular, for the years 2005 and 2006, posterior means for the latent Gaussian variable are high in the New Orleans area, which can be explained by high default rates following Hurricane Katrina in 2005. In 2008 and subsequent years, posterior means are high in the so-called “bubble” states California, Florida, Nevada, and Arizona, where house prices rose particularly rapidly in the run-up to the subprime mortgage crisis (Haughwout et al., 2011). Comparing posterior means of the tree-boosted spatio-temporal frailty model with those of the linear spatio-temporal model, we find similar patterns over space and time.
Figure 6: Posterior mean for the latent Gaussian process in the tree-boosted spatio-temporal frailty model when training on data up to the year 2021.
In Figure 7, we report the estimated covariance parameters for the different expanding window training data sets for the linear spatial, the linear spatio-temporal, and the tree-boosted spatio-temporal frailty models. For the latter, estimated variance and range parameters are relatively constant until the year 2020, when the variance parameter increases from 0.46 to 1.33, and in the following year 2021, the range parameter for time decreases from 2.94 to 1.05. These estimates are likely a consequence of the sudden atypical one-year default spike in 2020 during the COVID-19 pandemic. Since default mechanisms were likely very different in this year compared to the past due to an artificial external shock, less variation in the response variable is explained by the fixed effects and the latent Gaussian variable becomes more important as can be inferred from the higher marginal variance. Subsequently, defaults return to normal levels in 2021, and the estimated correlation with previous years is lower when including 2020 in the training data as can be seen from the lower estimated range parameter for time.
Figure 7: Estimated covariance parameters for different expanding window training data sets.
For understanding the function $F(\cdot)$ of the tree-boosted spatio-temporal frailty model, we use SHAP values and SHAP dependence plots (Lundberg and Lee, 2017). Specifically, we consider SHAP values for the model trained on data up to the year 2013. We chose this training window because for the corresponding prediction year 2014, there are pronounced differences in the accuracy of predictive default probabilities between the tree-boosted and linear spatio-temporal frailty models. SHAP values are calculated using 10’000 randomly selected instances of the training data. This subsampling is done to reduce the computational complexity and to avoid that SHAP dependence plots are overcrowded. Using a larger random subsample or a subsample chosen with a different random number generator seed gives almost identical results (results not reported). Figure 8 shows the SHAP values. The predictor variables are ordered according to the average of the absolute values of the SHAP values. According to these results, the eight most important predictor variables in descending order are the credit score (credit_score), the interest rate spread (ir_spread), the current loan-to-value (cnt_ltv), the indicator whether multiple borrowers are obligated to repay (multiple_borrowers), the age of the loan (n_months), the indicator whether the mortgage is used for purchase (loan_purpose = P), the original debt-to-income (orig_dti), and the original unpaid principal balance (orig_upb). For the six numeric variables thereof, we report the corresponding SHAP dependence plots in Figure 9. In all dependence plots, we observe strong interaction effects. This can be inferred from the large vertical scatter of the SHAP values and the systematic colored relationship with other predictor variables. For example, the slope of the SHAP values for the credit score variable (credit_score) is less negative if the interest rate spread variable (ir_spread) is positive. I.e., when interest rate spreads are high, the individual credit score matters less compared to when interest rate spreads are low. In general, the interaction effects are particularly pronounced for very large and small values at the boundaries of the predictor variables. Further, we find strong non-linear effects for the age of the loan (n_months) and the original unpaid principal balance (orig_upb). For instance, the age of a loan is positively related to the default probability up to a certain age of approximately three years, and then the effect flattens out and even starts to slightly decrease. In addition, the variables interest rate spread (ir_spread), current loan-to-value (cnt_ltv), and original debt-to-income (orig_dti) also have clearly non-linear relationships with slopes that change markedly for large and small values of these variables. For instance, increasing interest rate spreads are related to higher default probabilities, but this effect only holds up to a certain level of approximately three percent after which the relationship levels off and further increases in interest rate spreads are not associated with higher default probabilities. Furthermore, the debt-to-income (orig_dti) appears to have an approximately logistic-shaped effect on the default probability with the relationship being almost flat for both small debt-to-income values below 20 percent and large values above 50 percent, and in between, the effect of debt-to-income is approximately linear with a slope that interacts, among other things, with the loan age. We conclude that the presence of interaction and non-linear effects is likely the reason why the tree-boosted spatio-temporal frailty model outperforms the linear spatio-temporal model in terms of prediction accuracy.
Figure 8: SHAP values for the tree-boosted spatio-temporal frailty model when training on data up to the year 2013.
Figure 9: SHAP dependence plots for the tree-boosted spatio-temporal frailty model when training on data up to the year 2013.
4 Conclusion
We introduce a novel machine learning model combining tree-boosting with a latent spatio-temporal Gaussian process that accounts for frailty correlation. We compare our proposed model to an independent linear hazard model, a linear model with spatial frailty effects, and a linear model with spatio-temporal frailty effects and find that predictive default probabilities and predictive loan portfolio loss distributions obtained with our novel model are more accurate. Using interpretability tools for machine learning models, we provide evidence that the reasons for this better performance are interactions and non-linear effects in the predictor variables that cannot be captured by a linear functional form. In addition, we find that there are relatively strong spatio-temporal frailty effects.
Acknowledgments
We would like to thank Raffaella Calabrese for her helpful advice for gathering the mortgage credit risk data. This research was partially supported by the Swiss Innovation Agency - Innosuisse (grant number ‘55463.1 IP-ICT’).
Appendix A Appendix
A.1 Additional results for prediction of individual default probabilities
Figure 10: Temporal out-of-sample test H-measure (higher = better).
Figure 11: Temporal out-of-sample test Expected Calibration Error (ECE) (lower = better).
A.2 Additional results for prediction of loan portfolio loss distributions
Figure 12: Differences between means of the predictive loss distributions and the realized portfolio losses.
A.3 Posterior mean for the latent Gaussian process of the linear spatio-temporal model
Figure 13: Posterior mean for the latent Gaussian process in the linear spatio-temporal model when training on data up to the year 2021.
A.4 Tuning parameters
Tuning parameter Candidate values
Number of trees {1,2,…,1000}
Learning rate {10,1,0.1}
Maximal tree depth {2,3,5,10}
Minimal number of samples per leaf {10,100,1000}
L2 regularization {0,1,10}
Table 4: Candidate tuning parameters for the tree-boosted spatio-temporal frailty model.
Figure 14: Selected tuning parameters.
A.5 Summary statistics of predictor variables
Min. Q.25% Median Mean Q.75% Max.
credit_score 300.000 698.000 743.000 734.625 779.000 850.000
longitude -123.758 -104.952 -87.513 -92.403 -80.247 -67.866
latitude 25.531 35.211 39.520 38.657 41.968 48.428
insurance_percent 0.000 0.000 0.000 5.344 0.000 55.000
orig_dti 1.000 27.000 34.679 34.339 41.000 65.000
orig_cltv 2.000 66.000 80.000 75.735 88.000 534.000
orig_upb 9000.000 106000.000 160000.000 186612.656 243000.000 1581000.000
cnt_ltv 0.000 59.252 72.373 69.203 79.351 526.587
ir_spread -3.140 -0.010 0.610 0.693 1.265 7.080
n_months 0.000 12.000 29.000 39.967 57.000 274.000
Table 5: Summary statistics for the numeric predictor variables.
Level Count
occupancy I 167551
P 2008823
S 80154
nr_units 1 2181346
2 55477
3 10304
4 9401
loan_purpose C 583483
N 763290
P 909755
first_time_homebuyer 0 1972357
1 284171
msa 0 395311
1 1861217
multiple_borrowers 0 1027007
1 1229521
year 2000 21630
2001 48187
2002 53878
2003 52688
2004 41752
2005 53326
2006 71582
2007 92199
2008 112567
2009 128666
2010 125085
2011 119286
2012 117396
2013 105321
2014 101971
2015 112653
2016 118557
2017 122804
2018 132613
2019 145128
2020 149756
2021 121534
2022 107949
Table 6: Summary statistics for the categorical predictor variables.